OpenStudy (anonymous):

What is the square root of u - 15 = 7

OpenStudy (zzr0ck3r):

\(\sqrt{u-15}=7\) Square both sides \(\sqrt{u-15}^2=7^2=49\) giving \(u-15=49\)

OpenStudy (tkhunny):

It's an equation using the variable 'u'. What else would it be?